Homework: Section 3.1 Homework 3.1.32 (Type an integer or a decimal. Do not round.) Clear All
Answer #1

Given:

Mean height of 16 students in the statistics class to be 68 inches

Then Danielle walks in late, Danielle is 66.3 inches tall.

Find: Mean height of 17 students.

Therefore, Total height of 16 students is = (68*16) = 1088 inches

Therefore, Average height of 17 students is = (1088 + 66.3) / 17 = 67.9

The mean height of 17 students in the class is 67.9 inches.